OpenStudy (asylum15):

Evaluate the indefinite integral :

OpenStudy (asylum15):

\[\int\limits (x^3 + \frac{ 4 }{ x^2 } - e^{-2x}) dx\]

OpenStudy (asylum15):

Hi Chloro! :)

OpenStudy (asylum15):

Can I show you what I did?

OpenStudy (asylum15):

\[\frac{ x^4 }{ 4 }\]

OpenStudy (asylum15):

\[- \frac{ 4 }{ x }\]

OpenStudy (asylum15):

Can i ask, how does 4/x^2 become -4/x again?

OpenStudy (asylum15):

No,I followed a rule :D

OpenStudy (asylum15):

1/a^b = a^-b

OpenStudy (asylum15):

so 4x ^ -2?

OpenStudy (asylum15):

Ok, back to the equation :D

OpenStudy (asylum15):

-e^-2x = 1/2e ^-2x?

OpenStudy (asylum15):

\[e^{ax} = \frac{ 1 }{ a } e^{ax}\]

OpenStudy (asylum15):

?

OpenStudy (anonymous):

\[\frac{1}{2e^{2x}} + \frac{x^4}{4} - \frac{4}{x} + C\]

OpenStudy (anonymous):

you have all the right ideas, just put them together!

OpenStudy (asylum15):

How do i get 1/2e^-2x to 1/2e^2x mate?

OpenStudy (anonymous):

you know that \[e^{-2x} = \frac{1}{e^{2x}} \] right? so just integrate that as you would normally

OpenStudy (anonymous):

I think about it interms of the antiderivative, ie. what would i take the derivative of to get this.

OpenStudy (asylum15):

1/2 goes under the line and becomes 2 then?

OpenStudy (anonymous):

well you get 1/2 * 1/e^2x so kinda!

OpenStudy (asylum15):

Could I ask you a quick question?

OpenStudy (anonymous):

sure